OpenStudy (anonymous):

Let f''(x)=12x^2+6,f(1)=4,f'(1)=4. Then f(x) =

OpenStudy (anonymous):

6 x+4 x^3 + b = f'(x) f' (1) = 10 + b = 4 b = -6 f' (x) = 6 x+4 x^3 -6 f (x) = -6 x+3 x^2+x^4 + constant